xx PeriodicCarpet

PeriodicCarpet

Erik Schnetter <eschnetter@perimeterinstitute.ca>

January 09 2012

Abstract

1 Introduction

PeriodicCarpet applies periodic boundary conditions to Carpet grids. In particular, it provides correct behaviour for the case where a refinement level overlaps the periodic boundaries, and may therefore be made up of more than one component on the same process. Thorn Periodic, by constrast, will not handle this case properly.

2 Using This Thorn

The usage is rather simple and resembles that of Periodic, with boolean parameters periodic, periodicxx, periodicxy and periodicxz. An important point to notice is that if CarpetRegrid2 is used, its

parameters symmetryxperiodicxx, symmetryxperiodicxy and symmetryxperiodicxz have to be set

accordingly.

3 Parameters

11periodic 10Scope: privatex01  BOOLEAN

11Description: Periodic boundary conditions in all directions

11        10  x01  Default: no

11

11periodicxx 10Scope: privatex01  BOOLEAN

11Description: Periodic boundary conditions in x-direction

11        10  x01  Default: no

11

11periodicxy 10Scope: privatex01  BOOLEAN

11Description: Periodic boundary conditions in y-direction

11        10  x01  Default: no

11

11periodicxz10Scope: privatex01  BOOLEAN

11Description: Periodic boundary conditions in z-direction

11        10  x01  Default: no

11

11verbose 10Scope: privatex01  BOOLEAN

11Description: Produce screen output while applying boundary conditions

11        10  x01  Default: no

11

4 Interfaces

General Implements: periodiccarpet

Inherits: grid Uses header: carpet.hh mpixstring.hh dist.hh

Provides: PeriodicxApplyVI to PeriodicxApplyVN to PeriodicxApplyGI to PeriodicxApplyGN to

5 Schedule This section lists all the variables which are assigned storage by thorn Carpet/PeriodicCarpet. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxPARAMCHECK x periodiccarpetxparamcheck x register periodic boundary conditions x

11x11Language:11c
11 x 11Options: 11level
11 x 11Type: 11function

11

SymmetryRegister x periodiccarpetxregisterbc x register periodic boundary conditions x

11x11Language:11c
11 x 11Type: 11function

BoundaryConditions x periodiccarpetxapplybc x apply periodic boundary conditions x

11x11Language:11c
11 x 11Type: 11function